# Authorship Style Transfer with Policy Optimization

### v1

@@ Acknowledgements
This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200006. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ Appendix A More Experimental Results
STRAP LLaMA-7B - ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
STYLL GPT-3.5-turbo - ${}^{\ddagger}$${}^{\dagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
BLOOM-7B - ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
LLaMA-7B - 0.159 0.845 ${}^{\ddagger}$ ${}^{\ddagger}$ 0.59
SFT LLaMA-7B - ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Astrapop-PPO LLaMA-7B T + LP ${}^{\ddagger}$ ${}^{\ddagger}$ 0.767 ${}^{\ddagger}$ ${}^{\ddagger}$
A + LP ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
T + A + LP ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\dagger}$${}^{\dagger}$
Astrapop-DPO LLaMA-7B T + LP ${}^{\ddagger}$ ${}^{\ddagger}$ul0.761 ${}^{\dagger}$ ${}^{\ddagger}$
A + LP ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
@@
Astrapop-CPO LLaMA-7B T + LP ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\dagger}$ ${}^{\ddagger}$
A + LP ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
T + A + LP 0.165 ${}^{\ddagger}$ ${}^{\ddagger}$${}^{\dagger}$ ${}^{\dagger}$
@@
Zero-shot GPT-3.5-turbo - ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
LLaMA-7B - ${}^{\ddagger}$ ${}^{\dagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
STYLL BLOOM-7B - ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
LLaMA-7B - ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
STRAP / SFT LLaMA-7B - ${}^{\ddagger}$ ${}^{\ddagger}$ 0.917 ${}^{\ddagger}$ ${}^{\ddagger}$
STEER LLaMA-7B TSS + F + MS [^9] TSS, F, and MS are the three components of the reward function for STEER, which stand for target style strength, fluency, and meaning similarity, respectively. ${}^{\ddagger}$ 0.926 ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Astrapop-PPO LLaMA-7B T + LP ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
A + LP ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
T + A + LP ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Astrapop-DPO LLaMA-7B T + LP ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
A + LP ${}^{\ddagger}$ ${}^{\ddagger}$ul0.917 ${}^{\ddagger}$ ${}^{\ddagger}$
T + A + LP ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Astrapop-CPO LLaMA-7B T + LP 0.655 ${}^{\ddagger}$ ${}^{\ddagger}$ 0.827 0.662
A + LP ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
@@
A.1 More Baseline LLMs
@@
A.2 Reward Function Ablation Study
@@ B.1 Statistical Significance Test
We perform a resampled paired t-test on all results. Specifically, we randomly draw subsets from the test set and perform paired t-tests on the scores of the subsets. For the individual authorship style transfer task, we sample at the author level since the style model works at the author-level. For the community authorship style transfer task, we sample at the document level. The hyperparameters for the resampling t-test are shown in Table 6.
@@ B.2 Hyperparameters
Learning Rate KL coef / beta Batch Size
@@ B.3 Model Input Formats
Method Model
Prompt
Zero-shot LLaMA-7B
@@ B.4 Hardware and Runtime
We report the training hardward and runtime for all learning-based approaches in this work in Table 15.
@@
Table 15: Training hardward and runtime for the learning-based approaches. For DPO, CPO, and STEER QUARK, the two time numbers are the data generation time on a single A40 GPU and the training time, respectively. We do not report the runtime for the paraphrase data generation since its done through the OpenAI API.
@@ C.1 Use of Existing Artifacts
We list all existing artifacts we use in this work with their licenses and links in Table 16. The number of parameters of the models are shown in the same table in parentheses. The artifacts are under various licenses, but all permit the use for research purposes. All artifacts listed are allowed to be used in this work.

### v2

@@ Acknowledgements
This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200006, and in part by the Defense Advanced Research Projects Agency (DARPA) under Agreement No. HR00112490374. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, DARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ A.1 Full Automatic Evaluation Results
STRAP LLaMA-7B - 0.088 0.793 0.650 0.414 0.30
@@
BLOOM-7B - 0.117 0.796 0.546 0.408 0.37
LLaMA-7B - 0.159 0.845 0.529 0.440 0.59
SFT LLaMA-7B - 0.137 0.707 0.754 0.484 0.32
Astrapop-PPO LLaMA-7B $T+(LP^{\alpha}-1)$ 0.119 0.753 0.767 0.480 0.29
$A+(LP^{\alpha}-1)$ 0.111 0.761 0.710 0.454 0.37
[…]
$A+(LP^{\alpha}-1)$ 0.135 0.739 0.729 0.479 0.35
$T+A+(LP^{\alpha}-1)$ul0.164 0.748 0.733 0.507 0.44
Astrapop-CPO LLaMA-7B $T+(LP^{\alpha}-1)$ 0.151 0.743 0.749 0.501 0.38
$A+(LP^{\alpha}-1)$ 0.146 0.731 0.721 0.485 0.33
[…]
@@
Zero-shot GPT-3.5-turbo - 0.005 0.811 0.885 0.240 0.013
LLaMA-7B - 0.022 0.880 0.738 0.321 0.033
STYLL BLOOM-7B - 0.049 0.673 0.828 0.388 0.065
LLaMA-7B - 0.210 0.832 0.854 0.598 0.227
STRAP / SFT LLaMA-7B - 0.286 0.785 0.917 0.659 0.300
STEER LLaMA-7B $TSS+F+MS$ [^12] TSS, F, and MS are the three components of the reward function for STEER, which stand for target style strength, fluency, and meaning similarity, respectively. 0.334 0.926 0.879 0.699 0.348
Astrapop-PPO LLaMA-7B $T+(LP^{\alpha}-1)$ 0.299 0.800 0.905 0.665 0.313
$A+(LP^{\alpha}-1)$ 0.235 0.782 0.906 0.623 0.250
$T+A+(LP^{\alpha}-1)$ 0.240 0.788 0.908 0.628 0.256
Astrapop-DPO LLaMA-7B $T+(LP^{\alpha}-1)$ 0.490 0.843 0.915 0.767 0.499
[…]
$T+A+(LP^{\alpha}-1)$ 0.488 0.837 0.915 0.765 0.497
Astrapop-CPO LLaMA-7B $T+(LP^{\alpha}-1)$ 0.655 0.887 0.897 0.827 0.662
$A+(LP^{\alpha}-1)$ 0.456 0.835 0.909 0.749 0.467
@@
A.2 More Baseline LLMs
@@
A.3 Reward Function Ablation Study
@@
A.4 Human Evaluation
We conduct a human study on the community authorship transfer task. We randomly select 10 samples for each target native language from the test set for STYLL, STRAP, STEER, and Astrapop-CPO, and collect up to 3 annotations for each. The samples are evaluated in two dimensions, style confusion (SC) and content preservation (CP). The style confusion is a simpler version of the confusion score we use in the automatic evaluation. We show the annotators three examples each in the source style and the target style, and ask them “which is the style of the transferred text”. The confusion score is 1 if they select the target style, other with 0. We assess the content preservation using a 3-point Likert scale ranging from 0 to 2. The detailed instructions are shown in Table 6.
The results are shown in Table 5. Since the style classification task has been shown to be very difficult for humans Krishna et al. (2020); Hallinan et al. (2023), we perform an independent t-test on the results and find no statistically significant difference in style confusion in any model pairs. However, we observe statistically significant differences in content preservation, which indicates that both STRAP and STEER are significantly better than STYLL and Astrapop-CPO. Also, the results on content preservation are generally consistent with the SBERT in the automatic evaluation except for the STEER model.
SC CP
[…]
STEER 0.690 1.279
Astrapop-CPO 0.537 1.018
[…]
Style Confusion Question Based on the examples above, what is the style of the following text?
Content Preservation Similar Most of the meaning (75% or more) of the two passages is the same.
Somewhat Similar Large portions (50-75%) of the passages are the same, but there are significant sections that differ or are present in only one passage.
Not Similar Only small portions (less than 50%) of the passages are the same.
Question How similar are the following two texts?
Table 6: Instructions used in the human evaluation.
A.5 Case Study
Target Sample traviling is a very nice thing , it hepls you see new coultcurs and to meet new poeplo , there are alot of ways for traviling and i believe that the best way is to travil with a group led by a tour guide , [… more]
Source Text The influnce of advertisements on the customers is worth commendable. The advertisers are projecting thier goods to the customers in a 'larger than real' manner.' [… more]
Zero-shot Advertisements have a powerful influence on consumers, overstating the features and benefits of products to make them seem better than they actually are. [… more]
STYLL Advertisements are used to promote products, making them appear attractive and useful. Through exaggeration, advertisers present their products as having exceptional features. [… more]
STRAP The advertisements have a great effect on the customers and they should be praised. The advertisers make the products seem more better in the eyes of the customers rather than they really are. [… more]
STEER Advertising has a positive impact on customers, as it promotes products in a way that exaggerates their qualities. Advertisers often portray their products as superior to reality. [… more]
Astrapop-CPO because the advertisment have alot effec on the custumers , and make the products seem better than how they really are, [… more]
We show a transfer example on the community authorship transfer task in Table 7 for a simple qualitative case study. It shows that Astrapop-CPO successfully captures a common typo, “alot” and three main characteristics of the target style: using all lowercase, using space before comma, and high typo rate, while no other model is able to capture any of these.
@@ B.1 Statistical Significance Test
We perform a resampled paired t-test on all results. Specifically, we randomly draw subsets from the test set and perform paired t-tests on the scores of the subsets. For the individual authorship style transfer task, we sample at the author level since the style model works at the author level. For the community authorship style transfer task, we sample at the document level. The hyperparameters for the resampling t-test are shown in Table 8.
@@ B.2 Hyperparameters
Learning Rate KL coef / $\beta$ Batch Size
@@ B.3 Model Input Formats
Method Model Prompt
@@ B.4 Hardware and Runtime
We report the training hardware and runtime for all learning-based approaches in this work in Table 17.
@@
Table 17: Training hardware and runtime for the learning-based approaches. For DPO, CPO, and STEER QUARK, the two numbers for time are the data generation time on a single A40 GPU and the training time, respectively. We do not report the runtime for the paraphrase data generation since it is done through the OpenAI API.
B.5 Paraphraser Selection
We train a LLaMA paraphraser on GPT-3.5-turbo generated paraphraser data for two main reasons. First, the transfer pipeline is fully local with the LLaMA paraphraser, which is more cost-efficient and manageable. Second, more importantly, we find that the LLaMA paraphraser trained on the GPT-3.5-turbo generated data performs even better than GPT-3.5-turbo. Specifically, the trained LLaMA paraphraser achieves 0.764 SBERT cosine similarity on the test set, while GPT-3.5-turbo only has 0.738.
B.6 Reward Model Selection
We use LUAR for the experiments on the Reddit data since LUAR is the SOTA and most widely used authorship verification model on the Reddit dataset. However, preliminary experiments show that LUAR has only 0.53 accuracy on the ETS test set, while a trained classifier achieves 0.71 accuracy on the same test set. Therefore, we use the trained classifier as the reward model for the experiments on the ETS dataset. The reason for the large performance gap is that ETS has a countable number of classes (11 native languages) with plenty of training data to learn to accurately determine the author’s native language. Also, LUAR is a representation model that is designed to solve open-set problems in which the test data may have authors with textual styles never seen in the training collection, and this is in line with the setting of our low-resource transfer task on Reddit.
@@ C.1 Use of Existing Artifacts
We list all existing artifacts we use in this work with their licenses and links in Table 18. The numbers of parameters of the models are shown in the same table in parentheses. The artifacts are under various licenses, but all permit the use for research purposes. All artifacts listed are allowed to be used in this work.